OpenStudy (anonymous):

In tests i never understand this type of question - f(x) = 1-6x Solve f(x) = 0. Can someone explain how to solve this :S Cheers 8-)

OpenStudy (anonymous):

f(x)=0 gives u 1-6x=0 x=1/6

OpenStudy (angela210793):

Uzma...shouldn't f(x)=6x-1?

OpenStudy (anonymous):

is there - with f(x)? i wasn't sure its a break or negative

OpenStudy (angela210793):

tht's a minus a guess

OpenStudy (anonymous):

if negative, then u r right for sure :)

OpenStudy (angela210793):

:)

OpenStudy (angela210793):

f(x)=6x-1=0 6x=1 x=1/6 same answer though

OpenStudy (anonymous):

I dont think it was meant to be "negaative f"

OpenStudy (anonymous):

think it was just a dash,

OpenStudy (anonymous):

to say like "this is the question:" lol

OpenStudy (anonymous):

Neinn, it wasnt, Sorry (was just a dash), i get it now though (Y) thanks :D

OpenStudy (angela210793):

I think it is a minus

OpenStudy (angela210793):

anyway..he/she has both results whtever tht is,,,,,